OpenStudy (mtwizzle420):

What is the length of the hypotenuse of the right triangle shown below. @roz_mar23 the triangle will be in the comments

OpenStudy (mtwizzle420):



OpenStudy (anonymous):

\[A ^{2}+B ^{2}=C ^{2}\]

OpenStudy (anonymous):

\[11^{2}+14^{2}=\sqrt{317}\]

OpenStudy (mtwizzle420):

196+121=317

OpenStudy (anonymous):

\[\sqrt{317}=17.8\]

OpenStudy (mtwizzle420):

17.8?:)